OpenStudy (anonymous):

Find the ratio described as a fraction in simplest form. According to an organization's membership list, it has 1800 members who are male and 720 who are female. What is the ratio of members who are male to members who are female?

OpenStudy (anonymous):

the ratio is 5:2

OpenStudy (anonymous):

\[\frac{1800}{720}=\frac{5(\cancel{360})}{2(\cancel{360})}=\frac{5}{2}\]